# The Analysis of Willmore Surfaces and its Generalizations in Higher Dimensions

### v1

@@ II.1.4 Critical points of the Willmore energy under various constraints
Instead of looking at critical points of the Willmore functional among the set of all the immersions, one can restrict the set of admissible variations (for example, one need to fix some geometric constraints for the study of the Canham–Helfrich functional in cell biology). Three main possibilities are curently studied.
One possibility, is to study the minimizers (and more generally the critical points) of the Willmore functional under the constraints of fixing both the area of the immersed surface and the volume enclosed it encloses. Since the Willmore functional is scale-invariant, this question is equivalent to prescribing the isoperimetric ratio of the immersion $\vec{\Phi}\colon\Sigma\to\mathbb{R}^{3}$ given by
@@
$\beta^{\mathcal{I}}(R)\coloneqq\inf\left\{W(\vec{\Phi}):\vec{\Phi}\in\mathrm{Imm}(\Sigma;\mathbb{R}^{3}),\ \mathcal{I}(\vec{\Phi})=R\right\}.$
Keller–Mondino–Rivière [96] proved in 2014 that a minimizer of $\beta^{\mathcal{I}}(R)$ exists provided that this number verifies some strict inequality. This inequality has been proved for all $R\in\left((36\pi)^{\frac{1}{3}},+\infty\right)$ when $\Sigma=\mathbb{S}^{2}$ by Schygulla [179] in 2012, when $\Sigma=\mathbb{T}^{2}$ by Scharrer [174] in 2022 and by Ndiaye–Schätzle [153] in all the remaining cases in 2015.
@@
Another possibility, is to study the minimization of the Willmore energy under a constraint on the total mean curvature. Given an immersion $\vec{\Phi}\colon\Sigma\to\mathbb{R}^{3}$, we denote
$\mathcal{T}(\vec{\Phi})\coloneqq\frac{1}{vol_{g_{\vec{\Phi}}}(\Sigma)^{\frac{1}{2}}}\int_{\Sigma}H_{\vec{\Phi}}\,d\textup{vol}_{g_{\vec{\Phi}}}.$
By Hölder inequality, it holds
@@
$\beta^{\mathcal{T}}(R)\coloneqq\inf\left\{W(\vec{\Phi}):\vec{\Phi}\in\mathrm{Imm}(\Sigma;\mathbb{R}^{3}),\ \mathcal{T}(\vec{\Phi})=R\right\}.$
Scharrer–West [176] proved the existence of a minimizer for $R\in\left(0,\sqrt{2}\,\mathcal{T}(\mathbb{S}^{2})\right)\setminus\{\mathcal{T}(\mathbb{S}^{2})\}$ and studied the properties of the map $(R\in\mathbb{R}\mapsto\beta^{\mathcal{T}}(R))$. For the case $R=\mathcal{T}(\mathbb{S}^{2})$, they argue that the arguments developed in [96] do not apply.
@@
Is $\beta^{\mathcal{T}}(\mathcal{T}(\mathbb{S}^{2}))$ reached? Is there a minimizer for $\beta^{\mathcal{T}}(R)$ with $R\notin\left(0,\sqrt{2}\,\mathcal{T}(\mathbb{S}^{2})\right)\setminus\{\mathcal{T}(\mathbb{S}^{2})\}$?
@@
$\forall\vec{\Phi}\in\mathrm{Imm}(\Sigma;\mathbb{R}^{3}),\qquad W_{\bar{H}}(\vec{\Phi})\coloneqq\int_{\Sigma}(H_{\vec{\Phi}}-\bar{H})^{2}\,d\textup{vol}_{g_{\vec{\Phi}}}.$
The Helfrich flow has been introduced in 2006 [98], where Kohsaka–Nagasawa prove the short term existence of solution to the gradient flow of $W_{\bar{H}}$ with the enclosed volume and the area prescribed by some given constants $V_{0}$ and $A_{0}$. The cases when the initial data has small energy or when it is radially symmetric have been studied by Rupp–Scharrer–Schlierf [173]. This work extends earlier results of Rupp [171, 172] concerning the Willmore flow (i.e. the case $\bar{H}=0$) under a prescription of the isoperimetric ratio instead of a prescription of both the enclosed volume and the area. In order to study the flow for more general initial data, where singularities might appear, Scharrer–West [175] developed a bubbling analysis for constrained Willmore surfaces in the spirit of [19, 106].
@@
Let $\vec{\Phi}$ be a weak immersion of a surface $\Sigma$, as mentioned above it defines a smooth conformal class — i.e. a Riemann Surface — which is the one induced by the pull back metric metric $g_{\vec{\Phi}}=\vec{\Phi}^{\,\ast}g_{{\mathbb{R}}^{3}}$. The conformal class defined by this immersion is said to be a degenerate point of the map $\mathcal{C}$ assigning the conformal class of $(\Sigma,g_{\vec{\Phi}})$ in the Teichmüller space if and only if there exists a non zero holomorphic quadratic form [^4] An holomorphic quadratic form is an holomorphic section of the bundle given by $\wedge^{1,0}T^{\ast}\Sigma\otimes\wedge^{1,0}T^{\ast}\Sigma$ where $\Sigma$ is equipped with the conformal structure induced by $g_{\vec{\Phi}}=\vec{\Phi}^{\,\ast}g_{{\mathbb{R}}^{m}}$ $q$ such that
@@
$\vec{h}_{0}:=-\,\partial\pi_{\vec{n}}\,{\otimes}\,\partial\vec{\Phi},$
where $\pi_{\vec{n}}$ is the orthogonal projection linear map from ${\mathbb{R}}^{m}$ onto the normal space to the immersion $N_{x}\Sigma\coloneq(\vec{\Phi}_{\ast}T_{x}\Sigma)^{\perp}$ and $\cdot,\cdot>_{WP}$ is denoting the Weil–Peterson scalar product given in local conformal coordinates for which
@@
$\left<\phi\ dz\otimes dz,\psi\ dz\otimes dz\right>:=e^{-4\lambda}\ \Re\left[\overline{\phi}\,{\psi}\right]\ .$
Immersions in ${\mathbb{R}}^{3}$ which are degenerate points of the map $\mathcal{C}$} have been first studied in the late XIXth century under the name of isothermic surfaces (see [165]). There are the surfaces, such that, away from umbilic points there exists locally conformal coordinates which give the principal directions.
Assuming now that $(\Sigma,g_{\vec{\Phi}})$ is a non degenerate point under some constraint on the conformal class of the induced metric then there exists an holomorphic quadratic form $q$ such that
@@
The holomorphic quadratic form is a Lagrange multiplier [^5] The holomorphic quadratic form identifies to the tangent space to the moduli space (theorem 4.2.2 [93]) associated to the conformal class constraint. Weak immersions which are solving (II.2) are proved to be smooth ([18]).
@@
$\left\{\begin{array}[]{l}\displaystyle d\left(\vec{L}\cdot d\vec{\Phi}\right)=0\\[14.22636pt] \displaystyle d\left(\vec{L}\times d\vec{\Phi}+H\,d\vec{\Phi}\right)=0\end{array}\right$. (II.4)
coming respectively from Noether theorem for dilations and rotations which is itself equivalent to the so-called $(\vec{R},S)$ system (I.13) combined with the structure equation (I.14). For the higher dimensional counterparts to Willmore functional and in particular for the Dirichlet energy of the mean curvature of a weak immersion of 4-dimensional manifolds into ${\mathbb{R}}^{5}$, we are asking in Open Question VI.4 whether the conservation laws issued from Noether for dilations and rotations do correspond to the variation of this energy under some constraint.
@@ II.2 In higher dimensions
Even though, geometric properties of generalized Willmore functionals have already been the subject to lots of works, the study of the analytical properties is relatively new. We list below the few known results and propose different directions of research.
@@ II.2.1 Graham–Reichert functional
where $g_{\vec{\Phi}}\coloneqq\vec{\Phi}^{*}\xi$, $\xi$ is the flat metric on $\mathbb{R}^{5}$, $\mathrm{I\!I}_{\vec{\Phi}}$ is the second fundamental form of $\vec{\Phi}(\Sigma)$ and $H_{\vec{\Phi}}=\frac{1}{4}\mathrm{tr}_{g_{\vec{\Phi}}}\mathrm{I\!I}_{\vec{\Phi}}$ is its mean curvature. It turns out that the minus sign in the definition is sufficient to make the above functional unbounded from below, as proved in [74, 125]. The example constructed in [125] is based on the observation that the integrand of $\mathcal{E}_{GR}$ is a negative constant for $\mathbb{S}^{2}\times\mathbb{R}^{2}$, however this is not a closed submanifold. As mentioned by Graham in a private communication, a simple way to obtain a degenerating example of immersions from a closed manifold is to consider an ellipsoid with two semi-axis of length 1 and two others equal to some parameter $a\in(0,+\infty)$ that we can either send to 0 or to $\infty$ (since $\mathcal{E}_{GR}$ is scale invariant, both possibilities are equivalent and we recover $\mathbb{S}^{2}\times\mathbb{R}^{2}$ by sending $a\to+\infty$). At first glance, it seems that non-compact CMC immersions can provide an interesting source of examples for possible degeneracy of $\mathcal{E}_{GR}$ towards -$\infty$.
@@
Does there exist critical points of $\mathcal{E}_{GR}$ which are not conformal transformations of minimal hypersurfaces?
@@ II.2.2 Alternatives
The key idea of [23, 69], which has recently been developed in [6], is the link between $Q$-curvature and generalized Willmore energies. They show in [69], that one can define a notion of extrinsic GJMS operators on submanifolds. From these operators, one deduce a notion of extrinsic $Q$-curvature. In [23], the authors prove that for a 4-dimensional submanifold $\Sigma^{4}\subset M^{n}$ with $n\geq 5$, the integral of the difference between the extrinsic $Q$-curvature and the intrinsic $Q$-curvature of $\Sigma$ is a generalized Willmore energy. Hence, one could hope to find similarities between the study of the standard $Q$-curvature and generalized Willmore energies.
In order to study the space of immersions of a given hypersurface $\Sigma^{4}\subset\mathbb{R}^{5}$ into $\mathbb{R}^{5}$ and to provide a notion of ”best representation”, one would need a functional which is bounded from below and, as for the 2-dimensional case, would be a sort of conformally invariant distance of hypersurfaces to totally umbilic sets. One possibility has been introduced in [125] motivated by the generalization of the Dirichlet energy for the conformal Gauss map in dimension 2, see also [6] where Allen–Gover propose an alternative in any codimension using the normal tractor: if $\vec{\Phi}\colon\Sigma^{4}\to\mathbb{R}^{5}$ is a smooth immersion, then
@@
It is possible that $\mathcal{E}$ has a specific structure for which, there is no minimizers. In this case, it would be interesting to find another functional, coercive as well and which does possess minimizers. For instance, one can prove the following lower bound for $\mathcal{E}_{GR}$:
@@
II.2.3 Analysis in high dimension
In order to answer the previous questions, one would need to develop a refined analysis for submanifolds in the spirit of the approach presented here and most likely combined with the existing theory of varifolds. An important starting point already raised in [80, Remark 5.4.6], would be to generalize the study of Coulomb frames in higher dimensions but also to generalize the work of Müller–Šverák [151] on the construction of isothermal coordinates for immersed surfaces in a Euclidean space. The Coulomb frame are of major importance due to their relation to isothermal coordinates, see for instance Theorem IV.4 below or [80, Section 5.4]. In higher dimensions, isothermal coordinates do not exist in general, the Cotton and Weyl tensors are the obstruction. However, harmonic coordinates always exist on any manifold of any dimension $n\geq 2$ (isothermal coordinates are harmonic). In recent work [128, 129], the second and third authors show that for an immersion $\vec{\Phi}\colon\Sigma^{4}\to\mathbb{R}^{5}$, an estimate of the following form provides an atlas of harmonic coordinates on $\Sigma$ describing a $C^{1}$-differential structure with quantitative estimates on all the quantities involved

### v3

@@ II.1.4 Critical points of the Willmore energy under various constraints
Instead of looking at critical points of the Willmore functional among the set of all the immersions, one can restrict the set of admissible variations (for example, one needs to fix some geometric constraints for the study of the Canham–Helfrich functional in cell biology). Currently, three main possibilities are studied.
One possibility is to study the minimizers (and more generally the critical points) of the Willmore functional under the constraints of fixing both the area of the immersed surface and the volume it encloses. Since the Willmore functional is scale-invariant, this question is equivalent to prescribing the isoperimetric ratio of the immersion $\vec{\Phi}\colon\Sigma\to\mathbb{R}^{3}$ given by
@@
$\beta^{\mathcal{I}}(R)\coloneqq\inf\left\{W(\vec{\Phi}):\vec{\Phi}\in\mathrm{Imm}(\Sigma,\mathbb{R}^{3}),\ \mathcal{I}(\vec{\Phi})=R\right\}.$
Keller–Mondino–Rivière [97] proved in 2014 that there is a minimizer of $\beta^{\mathcal{I}}(R)$ provided that this number verifies some strict inequality. This inequality has been proved for all $R\in\left((36\pi)^{\frac{1}{3}},+\infty\right)$ when $\Sigma=\mathbb{S}^{2}$ by Schygulla [182] in 2012, when $\Sigma=\mathbb{T}^{2}$ by Scharrer [177] in 2022 and by Ndiaye–Schätzle [156] in all the remaining cases in 2015.
@@
Another possibility is to study the minimization of the Willmore energy under a constraint on the total mean curvature. Given an immersion $\vec{\Phi}\colon\Sigma\to\mathbb{R}^{3}$ (where $\Sigma$ is a Riemann surface, and thus carries an orientation), we denote
$\mathcal{T}(\vec{\Phi})\coloneqq\frac{1}{\textup{vol}_{g_{\vec{\Phi}}}(\Sigma)^{\frac{1}{2}}}\int_{\Sigma}H_{\vec{\Phi}}\,d\textup{vol}_{g_{\vec{\Phi}}}.$
By Hölder inequality, it holds that
@@
$\beta^{\mathcal{T}}(R)\coloneqq\inf\left\{W(\vec{\Phi})\colon\vec{\Phi}\in\mathrm{Imm}(\Sigma,\mathbb{R}^{3}),\ \mathcal{T}(\vec{\Phi})=R\right\}.$
Scharrer–West [179] proved the existence of a minimizer for $R\in\left(0,\sqrt{2}\,\mathcal{T}(\mathbb{S}^{2})\right)\setminus\{\mathcal{T}(\mathbb{S}^{2})\}$ and studied the properties of the map $(R\in\mathbb{R}\mapsto\beta^{\mathcal{T}}(R))$. For the case $R=\mathcal{T}(\mathbb{S}^{2})$, they proved that
$\beta^{\mathcal{T}}(\mathcal{T}(\mathbb{S}^{2}))=\inf\left\{W(\vec{\Phi}):\vec{\Phi}\in\mathrm{Imm}(\Sigma,\mathbb{R}^{3})\right\}.$
Hence, the existence of a minimizer for $\beta^{\mathcal{T}}(\mathcal{T}(\mathbb{S}^{2}))$ should be very related to Conjecture II.2.
@@
Is there a minimizer for $\beta^{\mathcal{T}}(R)$ with $R\notin\left(0,\sqrt{2}\,\mathcal{T}(\mathbb{S}^{2})\right)\setminus\{\mathcal{T}(\mathbb{S}^{2})\}$?
In the borderline case $R=0$, we know by continuity of the map $R\mapsto\beta^{\mathcal{T}}(R)$ that $\beta^{\mathcal{T}}(0)\leq 8\pi$. Knowing whether there exists a minimizer for $\beta^{\mathcal{T}}(0)$ thus amounts to knowing whether $\beta^{\mathcal{T}}(0)=8\pi$ is true (see [179]).
@@
$\forall\vec{\Phi}\in\mathrm{Imm}(\Sigma,\mathbb{R}^{3}),\qquad W_{\bar{H}}(\vec{\Phi})\coloneqq\int_{\Sigma}(H_{\vec{\Phi}}-\bar{H})^{2}\,d\textup{vol}_{g_{\vec{\Phi}}}.$
The Helfrich flow has been introduced in 2006 [99], where Kohsaka–Nagasawa proved the short-term existence of solutions to the gradient flow of $W_{\bar{H}}$ with the enclosed volume and the area prescribed by some given constants $V_{0}$ and $A_{0}$. The cases when the initial data has small energy or when it is radially symmetric have been studied by Rupp–Scharrer–Schlierf [176]. This work extends earlier results of Rupp [174, 175] concerning the Willmore flow (i.e. the case $\bar{H}=0$) under a prescription of the isoperimetric ratio instead of a prescription of both the enclosed volume and the area. In order to study the flow for more general initial data, where singularities might appear, Scharrer–West [178] developed a bubbling analysis for constrained Willmore surfaces in the spirit of [19, 109].
@@
Let $\vec{\Phi}$ be a weak immersion of a surface $\Sigma$, as mentioned above it defines a smooth conformal class — i.e. a Riemann surface — which is the one induced by the pull back metric $g_{\vec{\Phi}}=\vec{\Phi}^{\,\ast}g_{{\mathbb{R}}^{3}}$. The conformal class defined by this immersion is said to be a degenerate point of the map $\mathcal{C}$ assigning the conformal class of $(\Sigma,g_{\vec{\Phi}})$ in the Teichmüller space if and only if there exists a non zero holomorphic quadratic form [^6] A holomorphic quadratic form is a holomorphic section of the bundle given by $\wedge^{1,0}T^{\ast}\Sigma\otimes\wedge^{1,0}T^{\ast}\Sigma$ where $\Sigma$ is equipped with the conformal structure induced by $g_{\vec{\Phi}}=\vec{\Phi}^{\,\ast}g_{{\mathbb{R}}^{m}}$ $q$ such that
@@
$\vec{h}_{0}\coloneq-2\,\partial\pi_{\vec{n}}\,{\otimes}\,\partial\vec{\Phi},$
where $\pi_{\vec{n}}$ is the linear orthogonal projection map from ${\mathbb{R}}^{m}$ onto the normal space to the immersion $N_{x}\Sigma\coloneq(\vec{\Phi}_{\ast}T_{x}\Sigma)^{\perp}$ and $\cdot,\cdot>_{WP}$ denotes the Weil–Peterson scalar product given in local conformal coordinates for which
@@
$\left<\phi\ dz\otimes dz,\psi\ dz\otimes dz\right>\coloneq e^{-4\lambda}\ \Re\left[\overline{\phi}\,{\psi}\right]\ .$
Immersions in ${\mathbb{R}}^{3}$ which are degenerate points of the map $\mathcal{C}$} have been first studied in the late XIXth century under the name of isothermic surfaces (see [168]). These are the surfaces, such that, away from umbilic points there exists locally conformal coordinates which give the principal directions.
Assuming now that $(\Sigma,g_{\vec{\Phi}})$ is a non degenerate point under some constraint on the conformal class of the induced metric then there exists a holomorphic quadratic form $q$ such that
@@
The holomorphic quadratic form is a Lagrange multiplier [^7] The holomorphic quadratic form identifies to the tangent space to the moduli space (theorem 4.2.2 [94]) associated to the conformal class constraint. Weak immersions solving (II.2) are proved to be smooth ([18]).
@@
$\left\{\begin{array}[]{l}\displaystyle d\left(\vec{L}\cdot d\vec{\Phi}\right)=0,\\[14.22636pt] \displaystyle d\left(\vec{L}\times d\vec{\Phi}+H\,d\vec{\Phi}\right)=0,\end{array}\right$. (II.4)
coming respectively from Noether theorem for dilations and rotations which is itself equivalent to the so-called $(\vec{R},S)$ system (I.13) combined with the structure equation (I.14). For the higher dimensional counterparts to the Willmore functional and in particular for the Dirichlet energy of the mean curvature of a weak immersion of 4-dimensional manifolds into ${\mathbb{R}}^{5}$, we are asking in Open Question VI.4 whether the conservation laws issued from Noether for dilations and rotations do correspond to the variation of this energy under some constraint.
@@ II.2 In higher dimensions
Even though geometric properties of generalized Willmore functionals have already been the subject to a lot of works, the study of their analytical properties is relatively new. We list below the few known results and propose different directions of research.
@@ II.2.1 Graham–Reichert functional
where $g_{\vec{\Phi}}\coloneqq\vec{\Phi}^{*}\xi$, $\xi$ is the flat metric on $\mathbb{R}^{5}$, $\mathrm{I\!I}_{\vec{\Phi}}$ is the second fundamental form of $\vec{\Phi}(\Sigma)$ and $H_{\vec{\Phi}}=\frac{1}{4}\mathrm{tr}_{g_{\vec{\Phi}}}\mathrm{I\!I}_{\vec{\Phi}}$ is its mean curvature. It turns out that the minus sign in the definition is sufficient to make the above functional unbounded from below, as proved in [75, 129]. The example constructed in [129] is based on the observation that the integrand of $\mathcal{E}_{GR}$ is a negative constant for $\mathbb{S}^{2}\times\mathbb{R}^{2}$, however this is not a closed submanifold. As mentioned by Graham in a private communication, a simple way to obtain a degenerating example of immersions from a closed manifold is to consider an ellipsoid with two semi-axes of length 1 and two others equal to some parameter $a\in(0,+\infty)$ that we can either send to 0 or to $\infty$ (since $\mathcal{E}_{GR}$ is scale invariant, both possibilities are equivalent and we recover $\mathbb{S}^{2}\times\mathbb{R}^{2}$ by sending $a\to+\infty$). At first glance, it seems that non-compact CMC immersions can provide an interesting source of examples for possible degeneracy of $\mathcal{E}_{GR}$ towards -$\infty$.
@@
Are there critical points of $\mathcal{E}_{GR}$ which are not conformal transformations of minimal hypersurfaces?
@@ II.2.2 Alternatives
The key idea of [23, 69], which has recently been developed in [6], is the link between $Q$-curvature and generalized Willmore energies. They show in [69] that one can define an extrinsic version of the GJMS (Graham–Jenne–Mason–Sparling) operators [73] on submanifolds. From these operators, one deduces a notion of extrinsic $Q$-curvature. In [23], the authors prove that for a 4-dimensional submanifold $\Sigma^{4}\subset M^{n}$ with $n\geq 5$, the integral of the difference between the extrinsic $Q$-curvature and the intrinsic $Q$-curvature of $\Sigma$ is a generalized Willmore energy. Hence, one could hope to find similarities between the study of the standard $Q$-curvature and generalized Willmore energies.
In order to study the space of immersions of a given hypersurface $\Sigma^{4}\subset\mathbb{R}^{5}$ into $\mathbb{R}^{5}$ and to provide a notion of “best representation”, one would need a functional which is bounded from below and, as for the 2-dimensional case, would be a sort of conformally invariant distance of hypersurfaces to totally umbilic sets. One possibility has been introduced in [129] motivated by the generalization of the Dirichlet energy for the conformal Gauss map in dimension 2, see also [6] where Allen–Gover propose an alternative in any codimension using the normal tractor: if $\vec{\Phi}\colon\Sigma^{4}\to\mathbb{R}^{5}$ is a smooth immersion, then
@@
It is possible that $\mathcal{E}$ has a specific structure for which there are no minimizers. In this case, it would be interesting to find another functional, coercive as well and which does possess minimizers. For instance, one can prove the following lower bound for $\mathcal{E}_{GR}$:
@@
II.2.3 Analysis in high dimensions
In order to answer the previous questions, one would need to develop a refined analysis for submanifolds in the spirit of the approach presented here and most likely combined with the existing theory of varifolds. An important starting point already raised in [81, Remark 5.4.6], would be to generalize the study of Coulomb frames in higher dimensions but also to generalize the work of Müller–Šverák [154] on the construction of isothermal coordinates for immersed surfaces in a Euclidean space. The Coulomb frames are of major importance due to their relation to isothermal coordinates, see for instance Theorem IV.4 below or [81, Section 5.4]. In higher dimensions, isothermal coordinates do not exist in general, the Cotton and Weyl tensors are the obstruction. However, harmonic coordinates always exist on any manifold of any dimension $n\geq 2$ (isothermal coordinates are harmonic). In recent work [132, 133], the second and third authors show that for an immersion $\vec{\Phi}\colon\Sigma^{4}\to\mathbb{R}^{5}$, an estimate of the following form provides an atlas of harmonic coordinates on $\Sigma$ describing a $C^{1}$-differential structure with quantitative estimates on all the quantities involved